### v4

# Algebraic Gromov’s ellipticity of cubic hypersurfaces

Abstract.
We show that every smooth cubic hypersurface $X$ in ${\mathbb{P}}^{n+1}$, $n\geq 2$ is algebraically elliptic in Gromov’s sense. This gives the first examples of non-rational projective manifolds elliptic in Gromov’s sense. We also deduce that the punctured affine cone over $X$ is elliptic.
Key words and phrases:
spray, Gromov’s ellipticity, unirationality, stable rationality, cubic threefold, cubic hypersurface, affine cone.
2020 Mathematics Subject Classification. Primary 14J30, 14J70, 14M20; Secondary 14N25, 32Q56.
1. Introduction
Gromov’s ellipticity appeared (and was extremely useful) in complex analysis within the Oka-Grauert theory, see [Gro89], [For17a] and [For23]. In the same paper [Gro89] Gromov developed an algebraic version of this notion. We deal below exceptionally with algebraic ellipticity. Thus, all varieties and vector bundles in this paper are algebraic, and ‘ellipticity’ refers to algebraic ellipticity.
Let ${\mathbb{K}}$ be an algebraically closed field of characteristic zero and ${\mathbb{A}}^{n}$ resp. ${\mathbb{P}}^{n}$ be the affine resp. projective $n$-space over ${\mathbb{K}}$.
Given a smooth algebraic variety $X$, a spray of rank $r$ over $X$ is a triple $(E,p,s)$ where $p\colon E\to X$ is a vector bundle of rank $r$ with zero section $Z$ and $s\colon E\to X$ is a morphism such that $s|_{Z}=p|_{Z}$. A spray $(E,p,s)$ is dominating at $x\in X$ if the restriction $s|_{E_{x}}\colon E_{x}\to X$ to the fiber $E_{x}=p^{-1}(x)$ is dominant at the origin $0_{x}\in E_{x}\cap Z$ of the vector space $E_{x}$ i.e. $ds(T_{0}E_{x})=T_{x}X$.
One says that the variety $X$ is elliptic if it admits a spray $(E,p,s)$ which is dominating at each point $x\in X$, see [Gro89, 3.5A]. It is immediate from the definition that every elliptic manifold is unirational. An algebraic variety $X$ is said to be uniformly rational (or regular in the terminology of [Gro89, 3.5D]) if every point $x\in X$ has a neighborhood in $X$ isomorphic to an open subset of ${\mathbb{A}}^{n}$. It is called stably uniformly rational if $X\times{\mathbb{A}}^{k}$ is uniformly rational for some $k\geq 0$. A complete stably uniformly rational variety is elliptic, see [AKZ24, Theorem 1.3 and Corollary 3.7].
It is shown in [BB14, Example 2.4] that a rational smooth cubic hypersurface $X$ in ${\mathbb{P}}^{n+1}$, $n\geq 2$, is uniformly rational (cf. also [Gro89, Remark 3.5E′′′]). Hence $X$ is elliptic. There are examples of such hypersurfaces $X$ of any even dimension $n=2k\geq 2$.
As another example, consider a nodal cubic threefold $X$ in ${\mathbb{P}}^{4}$. It is well known that $X$ is rational. If $X$ has just a single node, say $P$, then the blowup ${\rm Bl}_{P}(X)$ of $X$ at the node is a smooth, uniformly rational (hence elliptic) threefold, see [BB14, Proposition 3.1]. The same holds for a small resolution $\tilde{X}\to X$ provided $X$ has several nodes and the resolution $\tilde{X}$ is an algebraic variety, see [BB14, Theorem 3.10]. Such an algebraic resolution exists, for instance, if $X$ has exactly 6 nodes, see [BB14, Proposition 3.8]. In the same spirit, one can construct examples of uniformly rational small resolutions of nodal quartic double solids, see [CZ24].
Yet another class of uniformly rational (and hence elliptic) projective varieties consists of smooth complete intersections of two quadrics in ${\mathbb{P}}^{n+2}$, $n\geq 3$, see [BB14, Example 2.5].
Gromov asked in [Gro89, 3.5B′′] whether any rational smooth projective variety is elliptic; the answer is still unknown. Gromov also discussed in [Gro89, 3.5B′′] a conjectural equivalence between ellipticity and unirationality. In the opposite direction, the question arises whether the ellipticity implies the (stable, uniform) rationality.
The answer to the latter question is negative. Indeed, for a certain natural number $n$ there exists a finite subgroup $F$ of $\operatorname{SL}(n,{\mathbb{C}})$ such that the quotient $Y=\operatorname{SL}(n,{\mathbb{C}})/F$ is stably irrational, see e.g. [Sal84, Theorem 3.6], [Bog87], [Pop11, Example 1.22] and [Pop13]. Being an affine homogeneous manifold of a semisimple group, $Y$ is flexible, that is, the subgroup $\operatorname{SAut}(Y)\subseteq\operatorname{Aut}(Y)$ generated by all one-parameter unipotent subgroups of $\operatorname{Aut}(Y)$ acts highly transitively on $Y$, see [AFKKZ13, Proposition 5.4]. The latter implies that $Y$ is elliptic, see e.g. [AFKKZ13, Proposition A.3]. Thus, $Y$ is an example of a stably irrational elliptic affine manifold.
Our aim is to find an irrational smooth projective variety $X$ which is Gromov’s elliptic. It is not certain that the above example can be explored for this purpose. Indeed, given a smooth elliptic affine variety $Y$, we lack the construction of a smooth elliptic completion $X$ of $Y$.
Nevertheless, we will establish the ellipticity of every smooth cubic threefold $X\subseteq{\mathbb{P}}^{4}$, see Theorem 1.1. By the celebrated Clemens-Griffiths theorem [CG72] such a threefold defined over ${\mathbb{C}}$ is irrational. According to Murre [Mur73] the same holds for every smooth cubic threefold defined over an algebraically closed field of characteristic different from 2.
Our main result is the following theorem.
Theorem 1.1.
A smooth cubic hypersurface $X\subseteq{\mathbb{P}}^{n+1}$ of dimension $n\geq 2$ is elliptic in the sense of Gromov.
Remarks 1.2.
1. It is not known whether there exists a stably (uniformly) rational smooth cubic threefold, see e.g. [CT19]. Also, no example of an irrational smooth cubic hypersurface of dimension $\geq 4$ is known. Recall that every smooth cubic hypersurface of dimension $\geq 2$ is unirational, see [Kol02] or [Huy23, Remark 5.13].
2. A smooth cubic threefold $X$ is far from being homogeneous. Indeed, due to Matsumura–Monsky’s theorem (see [MM63]) $\operatorname{Aut}(X)$ is a finite group.
3. Notice that any rational smooth projective surface $X$ admits a covering by open subsets isomorphic to ${\mathbb{A}}^{2}$. Hence it is uniformly rational and also elliptic. In the sequel we deal with cubic hypersurfaces of dimension $\geq 3$.
Recall the following open question.
Question 1.3 ([Gro89, Remark 3 3.5E′′′]).
Does the Gromov ellipticity of a smooth complete variety survive the birational maps generated by blowups and contractions with smooth centers?
There are some partial results concerning this question. The uniform rationality is preserved under blowups with smooth centers, see [BB14, Proposition 2.6] (cf. [Gro89, 3.5E-E′′]). Hence, also the ellipticity of a complete uniformly rational variety is preserved, see [AKZ24, Corollary 1.5]. For not necessary complete or uniformly rational elliptic varieties, the ellipticity is preserved under blowups with smooth centers under some additional assumptions, see [Gro89, Corollary 3.5.D′′], [LT17, Corollary 2], [KKT18] and [KZ23b, Theorem 0.1]. The preservation of ellipticity under blowdowns with smooth centers is unknown. See also the discussion in [Zai24].
Forstnerič [For17b] proved that every elliptic projective manifold $X$ of dimension $n$ admits a surjective morphism from ${\mathbb{C}}^{n}$, whose restriction to an open subset of ${\mathbb{C}}^{n}$ is smooth and also surjective. Due to Kusakabe [Kus22], this remains true for not necessary complete elliptic algebraic manifold $X$ over ${\mathbb{K}}$, provided one replaces ${\mathbb{C}}^{n}$ by ${\mathbb{A}}^{n+1}_{{\mathbb{K}}}$.
Recall that a generalized affine cone $\hat{Y}$ over a smooth projective variety $X$ defined by an ample ${\mathbb{Q}}$-divisor $D$ on $X$ is the affine variety
$\hat{Y}=\operatorname{Spec}\left(\bigoplus_{k=0}^{\infty}H^{0}\left(X,{\mathcal{O}}_{X}(\lfloor kD\rfloor)\right)\right),$
see, e.g., [KPZ13, Sec. 1.15]. In the case where $D$ is a hyperplane section of $X\subseteq{\mathbb{P}}^{n}$ the cone $\hat{Y}$ is the usual affine cone over $X$. The corresponding punctured cone $Y$ over $X$ is obtained from $\hat{Y}$ by removing the vertex of $\hat{Y}$; so $Y$ is a smooth quasiaffine variety.
Exploring the above results together with [KZ23b, Corollary 3.8 and Proposition 6.1] and Theorem 1.1, we deduce the following immediate corollary.
Corollary 1.4.
Let $X\subseteq{\mathbb{P}}^{n+1}$, $n\geq 2$ be a smooth cubic hypersurface and $Y$ be a punctured generalized affine cone over $X$ defined by an ample polarization of $X$. Then the following hold.
  • $Y$ is elliptic in Gromov’s sense.
  • There exist surjective morphisms ${\mathbb{A}}^{n+1}\to X$ resp. ${\mathbb{A}}^{n+2}\to Y$ which are smooth and surjective on appropriate open subsets $U\subseteq{\mathbb{A}}^{n+1}$ resp. $V\subseteq{\mathbb{A}}^{n+2}$.
  • If ${\mathbb{K}}={\mathbb{C}}$, then there exists a surjective morphism ${\mathbb{A}}^{n}\to X$ which is smooth and surjective on an appropriate open subset $U\subseteq{\mathbb{A}}^{n}$.
  • The monoid of endomorphisms $\operatorname{End}(Y)$ acts $m$-transitively on $Y$ for every natural number $m$.
2. Criteria of ellipticity
For the proofs of the following proposition see [Gro89, 3.5B], [For17a, Propositions 6.4.1 and 6.4.2], [LT17, Remark 3] and [KZ23b, Appendix B].
Proposition 2.1 (Gromov’s extension lemma).
Let $X$ be a smooth complete variety and $(E,p,s)$ be a spray on an open subset $U\subseteq X$ with values in $X$, that is, $p\colon E\to U$ is a vector bundle on $U$ with zero section $Z$ and $s\colon E\to X$ is a morphism such that $s|_{Z}=p|_{Z}$. Then for each $x\in U$ there is a smaller open neighborhood $V\subseteq U$ of $x$ in $X$ and a spray $(E^{\prime},p^{\prime},s^{\prime})$ on $X$ such that $(E^{\prime},p^{\prime},s^{\prime})|_{V}=(E,p,s)|_{V}$.
Recall that a variety $X$ is said to be locally elliptic if for any $x\in X$ there is a local spray $(E_{x},p_{x},s_{x})$ defined on a neighborhood $U$ of $x$ in $X$ and dominating at $x$ such that $s_{x}\colon E\to X$ takes values in $X$. The variety $X$ is called subelliptic if there is a family of sprays $\{(E_{i},p_{i},s_{i})\}_{i\in I}$ on $X$ which is dominating at each point $x\in X$, that is,
$T_{x}X={\rm span}({\rm d}s_{i}(T_{0_{i,x}}E_{i,x})\,|\,i\in I)\quad\forall x\in X.$
The following corollary on the equivalence of local and global ellipticity is given in [Gro89, 3.5B]; see also [KZ23a, Theorem 1.1 and Corollary 2.4].
Corollary 2.2 (Gromov’s Localization Lemma).
The ellipticity of a smooth algebraic variety $X$ is equivalent to its local ellipticity.
Sketch of the proof.
Obviously, ellipticity implies local ellipticity. Suppose that $X$ is locally elliptic. Choose a finite covering of $X$ that is equipped with sprays with values in $X$ dominating on the elements of this covering. We can assume that the vector bundles of these sprays are trivial. Thus, they can be decomposed into a direct sum of trivial line bundles. By Proposition 2.1 one can extend the resulting rank 1 sprays to sprays defined on the whole $X$, and then compose them. For rank 1 sprays the composition is again a spray, see [KZ23a, Proposition 2.1], which is dominating in our case. ∎
Remark 2.3.
The composition of sprays is defined in [Gro89, 1.3.B] for sprays of arbitrary ranks. It is mentioned in [Gro89, Section 1.3] that this does not give a spray, in general, if the ranks of participating sprays are $>1$. Note, however, that this circumstance is omitted in the proof of Lemma 3.5B in [Gro89].
Likewise, we obtain the following result.
Proposition 2.4 ([KZ23a, Theorem 1.1]).
The ellipticity of a smooth variety is equivalent to its subellipticity.
Given a spray $(E,p,s)$ with values in $X$ and a point $x\in X$, the constructible subset $O_{x}:=s(E_{x})\subseteq X$ is called the $s$-orbit of $x$. The proof of Proposition 2.4 leads to the following lemma, cf. [KZ23a].
Lemma 2.5.
A smooth variety $X$ of dimension $n$ is elliptic if for every $x\in X$ there exist local rank 1 sprays $(E_{i},p_{i},s_{i})$, $i=1,\ldots,n$ with values in $X$ defined on respective neighborhoods $U_{i}$ of $x$ such that the $s_{i}$-orbits $O_{i,x}$ of $x$ are curves with local parameterizations $s\colon(E_{i,x},0)\to(O_{i,x},x)$tale at $x$ whose tangent vectors at $x$ span $T_{x}X$.
Sketch of the proof.
Shrinking $U_{i}$ if necessary we may suppose that $(E_{i},p_{i},s_{i})$ is defined on the whole $X$ for $i=1,\ldots,n$, see Proposition 2.1. Since the tangent lines $ds(T_{0}E_{i,x})$ at $x$ to the $s_{i}$-orbits $O_{i,x}$, $i=1,\ldots,n$ span $T_{x}X$, the composition of these extended sprays is a spray of rank $n$ on $X$ dominating at $x$. This domination spreads to a neighborhood of $x$ in $X$. It follows that $X$ is locally elliptic (and also subelliptic). By Gromov’s Localization Lemma, see Corollary 2.2 (or alternatively by Proposition 2.4) $X$ is elliptic. ∎
In our proof of Theorem 1.1 in the next section we construct on any smooth cubic hypersurface $X$ of dimension $n\geq 3$ an $n$-tuple of independent rank 1 sprays $(E_{i},p_{i},s_{i})$ dominating in the orbit directions.
3. Ellipticity of cubic hypersurfaces
We use the following notation.
Notation 3.1.
For a subset $M\subseteq{\mathbb{P}}^{n}$ we let $\langle M\rangle$ be the smallest projective subspace which contains $M$. Given a smooth hypersurface $X\subseteq{\mathbb{P}}^{n+1}$ and a point $x\in X$ we let $S_{x}=\mathbb{T}_{x}X\cap X$, where $\mathbb{T}_{x}X$ stands for the projective tangent space of $X$ at $x$. Let ${\mathcal{C}}_{x}$ stand for the union of lines on $X$ through the point $x\in X$. Clearly, ${\mathcal{C}}_{x}\subseteq S_{x}$. Recall that for $n\geq 3$ every smooth cubic hypersurface $X\subseteq{\mathbb{P}}^{n+1}$ is covered by projective lines (this follows, for instance, from [CG72, Corollary 8.2]). Therefore, ${\mathcal{C}}_{x}$ has positive dimension for every $x\in X$.
Given a cubic threefold $X\subseteq{\mathbb{P}}^{4}$, the equality ${\mathcal{C}}_{x}=S_{x}$ holds if and only if $x$ is an Eckardt point of $X$. Recall that a point $x\in X$ is called an Eckardt point if there is an infinite number of lines on $X$ passing through $x$. In the latter case ${\mathcal{C}}_{x}$ is the cone with vertex $x$ over a plane elliptic cubic curve. Notice that a general cubic threefold has no Eckardt points, a cubic threefold can contain at most 30 Eckardt points (see [CG72, p. 315]), and the maximal number of 30 Eckardt points is attained only for the Fermat cubic threefold, see [CG72, p. 315], [Rou09] and [Huy23, Chapter 5, Remark 1.7].
Lemma 3.2.
Given a smooth cubic hypersurface $X\subseteq{\mathbb{P}}^{n+1}$ where $n\geq 3$, the tangent space $T_{x}X$ at a general point $x\in X$ is spanned by some $n$ tangent lines to projective lines on $X$ passing through $x$.
Proof.
Let us start with the case $n=3$, that is, let $X$ be a smooth cubic threefold. Through a general point $x\in X$ pass exactly 6 lines, see e.g. [AK77, Proposition (1.7)] or [Huy23, Chapter 5, Exercise 1.4]. More precisely, there is a proper closed subset $Y\subseteq X$ swept out by the “lines of the second type” on $X$, and through every point $x\in X\setminus Y$ pass exactly 6 distinct lines on $X$, see [Mur72, Lemma (1.19)]. These 6 lines form the cone ${\mathcal{C}}_{x}$. No four of them are coplanar, hence the tangent space $T_{x}X$ is spanned by the tangent lines to some three of these projective lines through $x$.
Let us show now that the latter property holds as well in higher dimensions. Given a smooth cubic hypersurface $X\subseteq{\mathbb{P}}^{n+1}$, $n\geq 4$ and a point $x\in X$, consider the projective variety
${\mathbb{P}}{\mathcal{C}}_{x}:=\{T_{x}l\,|\,x\in l\subseteq{\mathcal{C}}_{x}\}\subseteq{\mathbb{P}}T_{x}X$
where $l$ stands for a line on $X$. Clearly, $T_{x}X$ is spanned by the tangent lines $T_{x}l$ for $l\subseteq{\mathcal{C}}_{x}$ if and only if ${\mathbb{P}}{\mathcal{C}}_{x}$ is linearly nondegenerate in ${\mathbb{P}}T_{x}X\cong{\mathbb{P}}^{n-1}$. There exists an open dense subset $U\subseteq X$ such that for every $x\in U$ the projective cone ${\mathcal{C}}_{x}$ has codimension 2 in $\mathbb{T}_{x}X$. The dimension $\dim\,\langle{\mathbb{P}}{\mathcal{C}}_{x}\rangle$ is lower semicontinuous on $U$ and attains its maximal value, say $m$ on an open dense subset $U_{0}\subseteq U$. Let $\mathcal{V}\subseteq{\mathbb{P}}TX|_{U_{0}}$ be the subvariety swept out by the $\langle{\mathbb{P}}{\mathcal{C}}_{x}\rangle$ for $x\in U_{0}$.
Suppose that, contrary to the assertion of the lemma, $m<n-1$, so that $\mathcal{V}$ is a proper closed subset of ${\mathbb{P}}TX|_{U_{0}}$. Choose a general linear section $Y$ of $X$ by a subspace ${\mathbb{P}}^{4}\subseteq{\mathbb{P}}^{n+1}$. Then $Y$ is a smooth cubic threefold. Furthermore, for a general $x\in Y\cap U_{0}$ the proper subspaces ${\mathbb{P}}T_{x}Y\cong{\mathbb{P}}^{2}$ and $\langle{\mathbb{P}}{\mathcal{C}}_{x}\rangle\cong{\mathbb{P}}^{m}$ of ${\mathbb{P}}T_{x}X\cong{\mathbb{P}}^{n-1}$ are transversal. On the other hand, $T_{x}Y$ is spanned by the tangent lines to the lines on $Y$ passing through $x$. Thus, we have ${\mathbb{P}}T_{x}Y\subseteq\langle{\mathbb{P}}{\mathcal{C}}_{x}\rangle$. This contradiction ends the proof. ∎
Notation 3.3.
For a general point $u\in X$ we let $S^{*}_{u}$ be the set of points $x\in X$ such that $u\in\mathbb{T}_{x}X$. Let $x\in S^{*}_{u}$ be a general point and $u^{*}=\mathbb{T}_{u}X$ resp. $x^{*}=\mathbb{T}_{x}X$ be the corresponding points of the dual hypersurface $X^{*}\subseteq({\mathbb{P}}^{n+1})^{\vee}$. Then $u\in\mathbb{T}_{x}X$ if and only if $x^{*}\in\mathbb{T}_{u^{*}}X^{*}$, that is, $x\in S^{*}_{u}$ if and only if $x^{*}\in S_{u^{*}}$. It follows that $S^{*}_{u}$ is a hypersurface in $X$ passing through $u$. It is easily seen that ${\mathcal{C}}_{u}\subseteq S_{u}^{*}$, and so ${\mathcal{C}}_{u}\subseteq S_{u}\cap S_{u}^{*}$.
Lemma 3.4.
For a smooth cubic hypersurface $X\subseteq{\mathbb{P}}^{n+1}$, $n\geq 1$ one has $S_{u}\cap S_{u}^{*}={\mathcal{C}}_{u}$ for every $u\in X$.
Proof.
If $x\in S_{u}\cap S_{u}^{*}$ is different from $u$ then $l:=\langle x,u\rangle\subseteq\mathbb{T}_{x}X\cap\mathbb{T}_{u}X$. By the Bézout theorem, the bitangent line $l$ of $X$ is contained in $X$. It follows that $x\in{\mathcal{C}}_{u}$. Hence we have $S_{u}\cap S_{u}^{*}\subseteq{\mathcal{C}}_{u}$. Since also ${\mathcal{C}}_{u}\subseteq S_{u}\cap S_{u}^{*}$ one has $S_{u}\cap S_{u}^{*}={\mathcal{C}}_{u}$. ∎
Let again $X$ be a smooth cubic hypersurface in ${\mathbb{P}}^{n+1}$, $n\geq 1$. Following [BB14, Example 2.4] and [BKK13] let us consider the birational self-map $\tau_{u}\colon X\dasharrow X$ which sends a general point $x\in X$ to the third point $y=\tau(x)$ of intersection of the line $\langle x,u\rangle$ with $X$. Thus, $X$ cuts out on the line $\langle x,u\rangle$ the reduced divisor $x+u+y$. In fact, $\tau_{u}$ is well defined unless $x=u$ or $x\neq u$ and the line $\langle x,u\rangle$ is contained in $X$. Thus, $\tau_{u}\colon X\dasharrow X$ is regular on $X\setminus{\mathcal{C}}_{u}$.
Lemma 3.5.
Let $X\subseteq{\mathbb{P}}^{n+1}$, $n\geq 1$ be a smooth cubic hypersurface and $x,u\in X$ be such that $x\in X\setminus{\mathcal{C}}_{u}$. Then $\tau_{u}(x)=\tau_{u}(x^{\prime})$ for a point $x^{\prime}\in X$ if and only if either $u\in\mathbb{T}_{x}X$ (i.e. $x\in S_{u}^{*}$) and $x^{\prime}=x$, or $x\in\mathbb{T}_{u}X$ (i.e. $x\in S_{u}$) and $x^{\prime}=u$.
Proof.
Notice that by our assumption $x\neq u$ and $\langle x,u\rangle\not\subseteq X$. Suppose that $\tau_{u}(x)=\tau_{u}(x^{\prime})$ holds. Then we have $\langle x,u\rangle=\langle x^{\prime},u\rangle$. So the points $x,x^{\prime},u$ are alined. If $x^{\prime}\notin\{x,u\}$, then $\tau_{u}(x)=x^{\prime}$ and $\tau_{u}(x^{\prime})=x$. Applying $\tau_{u}$ once again we obtain $x^{\prime}=\tau_{u}(x)=\tau_{u}(x^{\prime})=x$, a contradiction. Thus, $x^{\prime}\in\{x,u\}$, and so the divisor cut out by $X$ on the line $\langle x,u\rangle$ is either $2x+u$, or $2u+x$. We have $x^{\prime}=x$ in the former case and $x^{\prime}=u$ in the latter case. Now the assertion follows. ∎
The following corollary is immediate.
Corollary 3.6.
  • (a)
    The morphism $\tau_{u}\colon X\setminus{\mathcal{C}}_{u}\to X$ contracts the hyperplane section $S_{u}\setminus{\mathcal{C}}_{u}$ to the point $u$ and fixes pointwise the variety $S_{u}^{*}\setminus S_{u}=S_{u}^{*}\setminus{\mathcal{C}}_{u}$ (see Lemma 3.4).
  • (b)
    The indeterminacy points of $\tau_{u}$ are contained in ${\mathcal{C}}_{u}$.
  • (c)
    The affine threefold $X\setminus S_{u}\subseteq X\setminus{\mathcal{C}}_{u}$ is invariant under $\tau_{u}$ and $\tau_{u}|_{X\setminus S_{u}}$ is a biregular involution with mirror $S_{u}^{*}\setminus{\mathcal{C}}_{u}$ that acts freely on $X\setminus(S_{u}\cup S_{u}^{*})$.
In the proof of Theorem 1.1 we use the following construction of a rank 1 spray on a smooth cubic hypersurface.
Proposition 3.7.
Let $X\subseteq{\mathbb{P}}^{n+1}$, $n\geq 3$ be a smooth cubic hypersurface, $y\in X$ be an arbitrary point and $x\in X$ be a general point. Let $u\in X\cap\langle x,y\rangle$ be a point different from $x$ and $y$. Choose a line $l\subseteq{\mathcal{C}}_{x}$ on $X$ through $x$, and let $C^{*}=\tau_{u}(l\setminus S_{u})\cong{\mathbb{A}}^{1}$. Then there exists a rank $1$ spray $(E,p,s)$ on $X$ such that the $s$-orbit $O_{y}$ of $y$ coincides with $C^{*}$ and $ds(T_{0}E_{y})$ is the tangent line to $O_{y}=C^{*}$ at $y$.
Proof.
Since $X$ cuts out on the line $\langle x,y\rangle$ the reduced divisor $x+u+y$ we have
$x,y\in U_{u}:=X\setminus(S_{u}\cup S_{u}^{*})\quad\text{and}\quad u,y\in U_{x}:=X\setminus(S_{x}\cup S_{x}^{*}),$
see Lemma 3.5. In particular, $x\notin\mathbb{T}_{u}X$. Hence the line $l\subseteq X$ passing through $x$ meets $\mathbb{T}_{u}X$ in a single point, say $z\in S_{u}=\mathbb{T}_{u}X\cap X$.
Fix an isomorphism $f\colon{\mathbb{P}}^{1}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}l$ that sends 00 to $x$ and $\infty$ to $z$. Then $f$ embeds ${\mathbb{A}}^{1}={\mathbb{P}}^{1}\setminus\{\infty\}$ onto $l\setminus\{z\}\subseteq X\setminus S_{u}$. Since $\tau_{u}|_{X\setminus S_{u}}\in\operatorname{Aut}(X\setminus S_{u})$, see Corollary 3.6(c), the map $\varphi_{u}:=\tau_{u}\circ f|_{{\mathbb{A}}^{1}}\colon{\mathbb{A}}^{1}\to C^{*}$ is an isomorphism.
By Corollary 3.6(c) $\tau_{x}|_{U_{x}}$ is a biregular involution acting freely on $U_{x}$ and interchanging $u$ and $y$. Since the projective line $\langle x,u\rangle$ is not tangent to $X$, we can choose an open neighborhood $V_{u}\subseteq U_{x}$ of $u$ such that also the line $\langle x,u^{\prime}\rangle$ is not tangent to $X$ for each $u^{\prime}\in V_{u}$. Letting $y^{\prime}=\tau_{u^{\prime}}(x)\in\langle x,u^{\prime}\rangle\cap X$ we have $\tau_{x}(u^{\prime})=y^{\prime}$. Letting $V_{y}=\tau_{x}(V_{u})\subseteq U_{x}$ the restriction $\tau_{x}|_{V_{u}}\colon V_{u}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}V_{y}$ is biregular, and so $V_{y}$ is a neighborhood of $y=\tau_{u}(x)$ in $X$.
Shrinking $V_{u}$ if necessary we may assume that for any $u^{\prime}\in V_{u}$ the tangent hyperplane $\mathbb{T}_{u^{\prime}}X$ does not pass through $x$, so that $z(u^{\prime}):=l\cap\mathbb{T}_{u^{\prime}}X\neq x$.
Consider the trivial ${\mathbb{P}}^{1}$-bundle $\pi\colon V_{u}\times l\to V_{u}$ over $V_{u}$ along with the constant section $Z_{0}=V_{u}\times\{x\}$ of $\pi$. The subset $\{z(u^{\prime})|u^{\prime}\in V_{u}\}\subseteq V_{u}\times l$ defines a section, say $Z^{\prime}$ of $\pi$ disjoint with $Z_{0}$. The restriction $\pi\colon(V_{u}\times l)\setminus Z^{\prime}\to V_{u}$ is a smooth ${\mathbb{A}}^{1}$-fibration with irreducible fibers.
There is an automorphism of the ${\mathbb{P}}^{1}$-bundle $\pi\colon V_{u}\times l\to V_{u}$ identical on the base $V_{u}$ that preserves $Z_{0}$ and sends $Z^{\prime}$ to a constant section, say, $Z_{\infty}$ disjoint with $Z_{0}$. This yields a trivialization
$F\colon V_{u}\times{\mathbb{A}}^{1}\stackrel{{\scriptstyle\cong|_{V_{u}}}}{{\longrightarrow}}(V_{u}\times l)\setminus Z^{\prime}$
that extends $f$ and sends the zero section of $V_{u}\times{\mathbb{A}}^{1}\to V_{u}$ to $Z_{0}$.
Consider the morphism
$\varphi\colon V_{u}\times{\mathbb{A}}^{1}\to X,\quad(u^{\prime},t)\mapsto\varphi_{u^{\prime}}(t):=\tau_{u^{\prime}}(F(t)).$
Let
$s=\varphi\circ(\tau_{x}\times\operatorname{id}_{{\mathbb{A}}^{1}})\colon V_{y}\times{\mathbb{A}}^{1}\to X,\,\,\,(y^{\prime},t)\mapsto\varphi_{u^{\prime}}(t)\,\,\text{where}\,\,u^{\prime}=\tau_{x}(y^{\prime})\in V_{u}.$
Consider also the trivial line bundle $p\colon E=V_{y}\times{\mathbb{A}}^{1}\to V_{y}$, where $p$ is the first projection, with zero section $Z=V_{y}\times\{0\}$. We have $s(y^{\prime},0)=y^{\prime}$, that is $s|_{Z}=p|_{Z}$. Thus, the triplet $(E,p,s)$ is a spray of rank 1 on $V_{y}$ with values in $X$. The $s$-orbit $O_{y}$ of $y$ coincides with $\varphi_{u}({\mathbb{A}}^{1})=C^{*}$ and the map $ds|_{T_{0}E_{y}}\colon T_{0}E_{y}\to T_{y}O_{y}$ is onto. Due to Proposition 2.1 $(E,p,s)$ can be extended to a spray on $X$. ∎
Remark 3.8.
We claim that the general $s$-orbits $O_{y^{\prime}}$ in $X$ are smooth affine conics whose closures are fibers of the conic bundle $\hat{X}\to{\mathbb{P}}^{2}$ resulting from the blowup of $\hat{X}\to X$ with center $l$. Indeed, let $y^{\prime}\in V_{y}$ be a general point. Then $\langle x,y^{\prime}\rangle$ being a general projective line in ${\mathbb{P}}^{4}$ through $y^{\prime}$, $X$ cuts the plane $L^{\prime}:=\langle l,y^{\prime}\rangle$ along the line $l$ passing through $x$ and the residual smooth conic $C_{y^{\prime}}$ passing through $u^{\prime}$ and $y^{\prime}$. So, the plane $L^{\prime}$ is $\tau_{u^{\prime}}$-invariant and $\tau_{u^{\prime}}$ interchanges $l$ and $C_{y^{\prime}}$ fixing their intersection points that are points of $l\cap S_{u^{\prime}}^{*}$, see Corollary 3.6(c). Furthermore, $C_{y^{\prime}}$ is the closure of the $s$-orbit $O_{y^{\prime}}=\tau_{u^{\prime}}(l\setminus\{z(u^{\prime})\})$.
We are now ready to prove Theorem 1.1.
Proof of Theorem 1.1.
Since every complete smooth rational surface is elliptic, see [KZ23b, Theorem 1.1], we may suppose that $n\geq 3$. Due to Corollary 2.2 it suffices to show that $X$ is locally elliptic. Fixing an arbitrary $y\in X$, choose a general point $x\in X$ and $n$ projective lines $l_{1},\ldots,l_{n}$ on $X$ through $x$ such that their tangent lines $T_{x}l_{i}$ span $T_{x}X$, see Lemma 3.2. Applying Proposition 3.7 to each of the lines $l_{i}$ yields rank 1 sprays $(E_{i},p_{i},s_{i})$, $i=1,\ldots,n$, on $X$ dominating at $y$ along their respective orbits $O_{i,y}$. We claim that the collection of sprays $\{(E_{i},p_{i},s_{i})\}_{i=1,\ldots,n}$ is dominating at $y$. The latter implies that $X$ is subelliptic, and so elliptic, see Corollary 2.2 (cf. also Lemma 2.5). The above domination is equivalent to the fact that the tangent lines $T_{y}O_{i,y}$ span the tangent space $T_{y}X$. By the construction of Proposition 3.7 we have
$T_{y}O_{i,y}=d\varphi_{i,u}(T_{x}l_{i})\,\,\,\text{where}\,\,\,\varphi_{i}=\tau_{u}\circ f_{i}\,\,\,\text{with}\,\,\,f_{i}\colon({\mathbb{A}}^{1},0)\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}(l_{i}\setminus\{z_{i}\},x).$
Therefore,
${\rm span}(T_{y}O_{1,y},\ldots,\,T_{y}O_{n,y})=d\tau_{u}({\rm span}(T_{x}l_{1},\ldots,\,T_{x}l_{n}))=d\tau_{u}(T_{x}X)=T_{y}X,$
which proves our claim. The domination at $y$ spreads to a neighborhood of $y$ in $X$, which gives the local ellipticity (and the subellipticity), hence also the ellipticity. ∎
4. An alternative proof
We suggest here an alternative proof of Theorem 1.1. The geometrical arguments used in Section 3 are now replaced by references to Forstnerič-Kusakabe’s theorem.
Lemma 4.1.
Let $X$ be a variety and $S\subseteq X\times{\mathbb{P}}^{n}$ be a subvariety of codimension at least $2$. Then for a general point $(x,a)\in X\times{\mathbb{P}}^{n}$ and a general projective line $L\subseteq{\mathbb{P}}^{n}$ through $a$ there exists a neighborhood $U_{x}$ of $x$ in $X$ such that $(U_{x}\times L)\cap S=\emptyset$.
Proof.
It suffices to show that ($(\{x\}\times L)\cap S=\emptyset$ for a general choice of $(x,a)\in X\times{\mathbb{P}}^{n}$ and $L$ in the Grassmannian ${\bf G}_{a}(1,n)$ of lines in ${\mathbb{P}}^{n}$ passing through $a$. Since $(x,a)\in X\times{\mathbb{P}}^{n}$ is general we have $(x,a)\notin S$. Let $L_{v}$ be a line passing through $a$ in direction of a general vector $v\in{\mathbb{P}}T_{a}{\mathbb{P}}^{n}\simeq{\mathbb{P}}^{n-1}$. Fixing $a$ assume to the contrary that ($(\{x\}\times L_{v})\cap S\neq\emptyset$ for general $x\in X$ and $v\in{\mathbb{P}}^{n-1}$. Then $\dim(S)\geq\dim(X)+n-1=\dim(X\times{\mathbb{P}}^{n})-1$. This contradicts the assumption that ${\rm codim}_{X\times{\mathbb{P}}^{n}}(S)\geq 2$. ∎
Corollary 4.2.
Let $X$ be a normal algebraic variety, $\psi\colon X\times{\mathbb{A}}^{n}\dashrightarrow X$ be a dominant rational map, and $S$ be the set of indeterminacy points of $\psi$. Then for a general point $(u,a)\in X\times{\mathbb{A}}^{n}$ and a general affine line $l\subseteq{\mathbb{A}}^{n}$ passing through $a$ there is a neighborhood $U_{u}$ of $u$ in $X$ such that $(U_{u}\times l)\cap S=\emptyset$, and so $\psi|_{U_{u}\times l}\colon U_{u}\times l\to X$ is a morphism.
Proof.
It suffices to extend $\psi$ to a rational map $\bar{\psi}\colon X\times{\mathbb{P}}^{n}\dasharrow X$, to replace $S$ by the indeterminacy set of $\bar{\psi}$ and to apply Lemma 4.1 in this new framework. ∎
The next proposition is an analog of Proposition 3.7.
Proposition 4.3.
Let $X$ be a smooth cubic hypersurface in ${\mathbb{P}}^{n+1}$ and $y$ be a point in $X$. Then there is a neighborhood $U_{y}$ of $y$ in $X$ and a rank $1$spray $(E,p,s)$ on $U_{y}$ with values in $X$ such that $E=U_{y}\times{\mathbb{A}}^{1}$, $p\colon U_{y}\times{\mathbb{A}}^{1}\to U_{y}$ is the first projection and a morphism $s\colon U_{y}\times{\mathbb{A}}^{1}\to X$ verifies the following conditions:
  • $s|_{Z}=p|_{Z}$ where $Z=U_{y}\times\{0\}$ is the zero section of $p\colon E\to U_{y}$;
  • the orbit map $s|_{E_{y}}\colon E_{y}\to s(E_{y})$ is smooth at the origin $0_{y}\in E_{y}$;
  • $ds|_{T_{0_{y}}E_{y}}$ sends a tangent vector to $E_{y}$ at $0_{y}$ to a general vector in $T_{y}X$.
Proof.
By Kusakabe’s theorem, see [Kus22], there is a surjective morphism $f:{\mathbb{A}}^{n+1}\to X$. Let $g=(\operatorname{id}_{X},f)\colon X\times{\mathbb{A}}^{n+1}\to X\times X$. Define a rational map $\tau\colon X\times X\dashrightarrow X$ by letting $\tau(u,x)=\tau_{u}(x)$. Then the composition
$\psi\colon X\times{\mathbb{A}}^{n+1}\stackrel{{\scriptstyle g}}{{\longrightarrow}}X\times X\stackrel{{\scriptstyle\tau}}{{\dashrightarrow}}X,\qquad(u,a)\mapsto\tau_{u}(f(a)),$
is a dominant rational map. Letting $I(\tau)\subseteq X\times X$ and $S:=I(\psi)\subseteq X\times{\mathbb{A}}^{n+1}$ be the indeterminacy sets of $\tau$ and $\psi$, respectively, we have $S\subseteq g^{-1}(I(\tau))$ and ${\rm codim}_{X\times{\mathbb{A}}^{n+1}}(S)\geq 2$.
Let now $L_{y}$ be a general line in ${\mathbb{P}}^{n+1}$ through $y$. It meets $X$ at general points $x$ and $u$ different from $y$. Recall that the indeterminacy set $I(\tau_{u})\subseteq X$ is contained in the union ${\mathcal{C}}_{u}$ of lines on $X$ passing through $u$, see the paragraph preceding Lemma 3.5. Since $L_{y}\not\subseteq X$, we have $x\notin I(\tau_{u})$, and so $(u,x)\notin I(\tau)$. Since $S\subseteq g^{-1}(I(\tau))$, for $a\in f^{-1}(x)$ we have $(u,a)\notin S$ and $\psi(u,a)=y$.
Fix a general affine line $l\subseteq{\mathbb{A}}^{n+1}$ through $a$. Let $U_{u}$ be a neighborhood of $u$ in $X$ such that $\psi$ is regular on $U_{u}\times l\simeq U_{u}\times{\mathbb{A}}^{1}$, see Corollary 4.2. Then $U_{y}:=\tau(U_{u}\times\{x\})$ is a neighborhood of $y$ in $X$ and $\tau_{x}\colon U_{u}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}U_{y}$ is an isomorphism. Composing $\psi$ with $(\tau_{x},\operatorname{id})$ one gets a morphism $s\colon U_{y}\times{\mathbb{A}}^{1}\to X$. This defines a desired rank 1 spray $(E,p,s)$ on $U_{y}$ with values in $X$. Indeed, $s(x,a)=y$ for any $a\in f^{-1}(x)$. Furthermore, $s|_{E_{y}}=s|_{(x,l)}=\tau_{u}(f(l))$ is the $s$-orbit of $y$. Since $x\in X$ is a general point, it is not a critical value of $f$. Hence for $a\in f^{-1}(x)$, the morphism $g=(\operatorname{id}_{X},f)\colon X\times{\mathbb{A}}^{n+1}\to X\times X$ is dominant at $(u,a)$. Since $l\subseteq{\mathbb{A}}^{n+1}$ is a general line passing through $a$, it follows that the differential $ds$ sends $T_{0_{y}}E_{y}$ to a general vector line in $T_{y}X$. ∎
Due to Gromov’s Localization Lemma (see Corollary 2.2) to prove Theorem 1.1 it remains to apply the following corollary.
Corollary 4.4.
$X$ is locally elliptic.
Proof.
Choose independent general vectors $v_{1},\ldots,v_{n}$ in $T_{y}X$ and the corresponding local rank 1 sprays $(E_{i},p_{i},s_{i})$ defined on a common neighborhood $U$ of $y$ in $X$ with values in $X$, see Proposition 4.3. Extend these sprays to sprays on the whole $X$, see Proposition 2.1. The resulting collection of sprays is dominating on $U$. This guarantees the subellipticity of $X$ on $U$. Now the claim follows by Proposition 2.4. ∎
Acknowledgments. It is our pleasure to thank Yuri Prokhorov for valuable discussions, references and comments on a preliminary version of this paper.